# Hardware-Accelerated Ray Tracing for Discrete and Continuous Collision Detection on GPUs

### v1

@@ IV-D Continuous Collision Detection Evaluation
For CCD, we benchmarked our piecewise-linear and quadratic B-spline sphere swept volume RT-CCD against discretized mesh pose trajectories checked with our Two-Way RT-DCD and cuRobo, the only other GPU-based library implementing CCD. Due to cuRobo’s sphere-based CA-style algorithm, its swept volumes are identical to our piecewise linear curves. Table I shows detection accuracy results, and Fig. 5 shows speed for different trajectory batch sizes and trajectory control points. We again omit our cubic B-spline results (see Sec. IV-B). Our piecewise-linear and discretized methods outpaced cuRobo across the board, with discretization being the fastest on dense scenes at the cost of recall rate. Quadratic B-splines had the best accuracy at slower speeds, but was still faster than cuRobo at most batch sizes.
@@ V-A1 RT-DCD Accuracy
As mentioned, our two-way volume mesh-to-mesh RT-DCD catches all penetrating collisions. However, it cannot detect mere surface contact between meshes as NVIDIA’s RT cores do not guarantee catching ray-triangle intersections where the ray is coplanar with the triangle or starts/ends at the triangle surface (such intersections have little use in graphics rendering). Mochi [8] quadruples the number of triangles to catch more of such collisions, but it still cannot catch all of them (such as triangles or edges fully contained in a coplanar triangle) without tracing longer rays with more complex post-processing logic. Since exact contact detection is often not important in collision-avoidance, we are content instead to detect only penetrating collisions with a much simpler approach.
@@ V-B Performance of RT-CCD
First, ray tracing against curves, particularly high-order curves, is much more expensive than against triangles as the narrow-phase intersection is not hardware-accelerated, running instead on normal CUDA cores. Thus it only benefits from hardware-acceleration during BVH traversal. Second, our RT-CCD requires expensive curve BVHs builds from scratch on the critical path, unlike RT-DCD which can reuse precomputed mesh BVHs in an IAS. Finally, curve BVHs are less efficient than for triangle meshes as the sphere extrusions lead to large AABBs, the bounding volume used in current NVIDIA RT cores. This is often addressed by splitting long curves into smaller segments or across multiple smaller AABBs, but this leads to longer BVH build times on our critical path which we found overall gives no speedup.

### v2

@@ IV-D Continuous Collision Detection Evaluation
For CCD, we benchmarked our piecewise-linear and quadratic B-spline sphere swept volume RT-CCD against discretized mesh pose trajectories checked with our two-way RT-DCD and cuRobo, the only other GPU-based library implementing CCD. Due to cuRobo’s sphere-based CA-style algorithm, its swept volumes are identical to our piecewise linear curves. Table I shows detection accuracy results, and Fig. 5 shows speed for different trajectory batch sizes and trajectory control points. We again omit our cubic B-spline results (see Sec. IV-B). Our piecewise-linear and discretized methods outpaced cuRobo on all except large batch sizes on simple scenes, with discretization being the fastest on dense scenes at the cost of recall rate. Quadratic B-splines had the best accuracy at slower speeds, but were still faster than cuRobo at most batch sizes for medium/dense scenes.
@@ V-A1 RT-DCD Accuracy
As mentioned, our two-way volume mesh-to-mesh RT-DCD catches all penetrating collisions. However, it cannot detect mere surface contact between meshes as NVIDIA’s RT cores do not guarantee catching ray-triangle intersections where the ray is coplanar with the triangle or starts/ends at the triangle surface (such intersections have little use in graphics rendering). Mochi [9] quadruples the number of triangles to catch more of such collisions, but it still cannot catch all of them (such as triangles or edges fully contained in a coplanar triangle) without tracing longer rays with more complex post-processing logic. Since surface contact with zero penetration can often be omitted in collision-avoidance (often subject in any case to numerical precision issues that one can avoid with slight robot or obstacle inflation), we are content instead to detect only penetrating collisions with a much simpler approach.
@@ V-B Performance of RT-CCD
First, ray tracing against curves, particularly high-order curves, is much more expensive than against triangles as up through the Ada generation GPUs used in this work, the narrow-phase intersection is not hardware-accelerated, running instead on normal CUDA cores. Thus it only benefits from hardware-acceleration during BVH traversal. This is expected to improve in Blackwell generation GPUs, which hardware-accelerate narrow-phase ray-traces against piecewise-linear curves and spheres.
Second, our RT-CCD requires expensive curve BVHs builds from scratch on the critical path, unlike RT-DCD which can reuse precomputed mesh BVHs in an IAS. Finally, curve BVHs are less efficient than for triangle meshes as the sphere extrusions lead to large AABBs, the bounding volume used in current NVIDIA RT cores. This is often addressed by splitting long curves into smaller segments or across multiple smaller AABBs, but this leads to longer BVH build times on our critical path which we found overall gives no speedup.
@@ VI Conclusion
Appendix A Further Implementation Details
We made a number of design decisions as a result of exploration and experimentation to build performant pipelines for collision detection with OptiX and NVIDIA ray-tracing cores. Key algorithm decision points included the following:
  • Using OptiX geometric acceleration structures (GAS) versus using an instance acceleration structure (IAS) for meshes/curves collision-checked in the same batch
  • Using versus omitting a custom single-level broad-phase elimination of inputs into the ray-tracing pipeline (either rays or collision geometries)
  • Choosing sub-batch sizes and the number of asynchronous CUDA streams algorithm sections run on
  • Optimizing acceleration structure builds for fast builds versus optimizing for fast ray traces against them
Since some of these choices are coupled, we will first introduce each concept and then list our recommendations and justifications for relevant collision detection algorithms.
A-A Decision Points and Tradeoffs
A-A1 GAS and IAS
In NVIDIA’s OptiX ray-tracing engine, geometric acceleration structures (GAS) and instance acceleration structures (IAS) are two types of acceleration structures which represent geometries to traced rays against (as of this writing, NVIDIA implements these acceleration structures as BVHs using AABB bounding volumes). A GAS consists of only geometric primitives (triangles, spheres, curves, etc.). on the other hand, an IAS is a tree of other acceleration structures which could be GASs or other IASs, each of which has an $SE(3)$ transform relative to its parent and can be included more than once. When geometries are duplicated, an IAS can be more efficient than a GAS, both for build times and memory usage. For example, if a mesh is duplicated across a scene, one can build a single GAS for the mesh and include it multiple times in an IAS in different poses. However, a single large GAS is more free to optimize a single acceleration structure for all the low-level primitives in a scene, which can sometimes lead to faster ray traces.
When tracing against a geometry such as a mesh without duplicated submeshes, there are differences between representing it as a pure GAS vs. a GAS in an IAS when the geometry is given a pose. With a pure GAS, the rays must be manually posed, for example in a ray generation program which runs on CUDA cores. With a posed GAS in an IAS, the pose is applied to the rays during the scene traversal. Thus we must carefully decide whether to represent the obstacles, batch-posed robots, and trajectories as GASs or IASs to achieve the best performance.
A-A2 Custom Single-Level Broad-Phase Elimination
Although full BVH construction and traversal are complex to implement with high performance on the GPU, a single-level custom broad-phase elimination pass can be very simple while producing significant reductions in inputs to the ray-tracing steps for collision detection. For example, in our experiments, often only $\sim$1/4th of single-level oriented bounding boxes (OBBs) encompassing robot links were in collision with OBBs representing the obstacles.
While the OptiX acceleration structures reduce the number of geometries each ray needs to be traced against, we can use our custom single-level broad-phase to reduce the number of rays traced at all along both the robot and obstacles, which also has a large impact on performance. We can also build simpler acceleration structures containing only robot links which are true collision candidates. However, since the correct links cannot be known ahead of time, this requires building new acceleration structures online rather than performing a fast in-place dynamic update on a pre-built acceleration structure online.
A-A3 Sub-Batch Size and Stream Count
Given a large batch of collision detection queries, we can process the whole batch together or split it into smaller sub-batches which are each checked using a single ray-tracing call and a single call of other pre- or post-processing CUDA kernels. For example, in DCD ObsToRobot and in our CCD we trace obstacle rays in one large batch, regardless of the number of poses or trajectories checked in the batch.
Increasing the size of one large batch can increase parallelism up to the point when regions of the GPU become saturated and produce bottlenecks. Bottlenecks can occur in different GPU regions, such as in compute (e.g., CUDA cores or RT cores) or in memory, hitting memory access or data bus throughput limits. Since these regions can often perform simultaneous operations with each other, we can sometimes relieve bottlenecks by splitting a large batch into smaller sub-batches. We can then schedule each sub-batch’s operations on a separate CUDA stream (a queue of sequential tasks which can execute in parallel with tasks queued on other streams). This provides opportunity for parts of a sub-batch’s compute pipeline with heavy use of some GPU region to execute in parallel with a different part of another sub-batch’s compute pipeline which doesn’t use the region as heavily.
On the other hand, too many small batches on too many streams may incur too much kernel launch and scheduling overhead. Thus, to handle large collision detection batch requests, we could use CUDA streams to divide the batch into sub-batches to maximize device occupancy, but the optimal sub-batch size depends on the algorithm and the particular GPU being used.
A-A4 Acceleration Structure Build Flags
OptiX provides different build options for acceleration structures depending on user priorities. Two of the main options switch between preferring a fast acceleration structure builds versus preferring faster ray traces against the resulting acceleration structures. In most cases, we chose faster ray traces, but in CCD, we must build curve GASs online, making fast build times more important. In particular, for piecewise linear curves, we found prioritizing fast builds resulted in better overall algorithm speeds.
Figure 6: Sphere representation of the Franka Panda robot used in our experiments.
A-B Choices for Algorithms
Based on experimentation, we settled on the following configurations. Note that these may be subject to change on different GPUs and on different variations of our algorithms.
A-B1 DCD ObsToRobot
When shooting rays from obstacles to robots, we use an Instance Acceleration Structure (IAS) to represent batched robot links across different robot poses. The IAS saves memory by avoiding duplication of link meshes, which would be needed by a Geometry Acceleration Structure (GAS). We use broad-phase collision detection to filter out any links that can never collide. We then gather selected link instances through stream compaction [26] and use them to build an IAS. On the other side, we found it was rare that obstacles are collision-free across all poses in a batch, so we do not filter them using the broad-phase collision detection results. In our testing, we found it was best to use 4 streams and 1024 poses per batch, which achieves high device occupancy and detection throughput.
(a) Dense Scene
(b) Medium Scene
(c) Simple Scene
Figure 7: Runtimes for discrete collision detection (DCD) algorithms in different scenes with cuRobo CUDA graphs disabled.
(a) Dense Scene
(b) Medium Scene
(c) Simple Scene
Figure 8: Runtimes for continuous collision detection (CCD) algorithms in different scenes with cuRobo CUDA graphs disabled.
A-B2 DCD RobotToObs
For this algorithm, we represent the obstacles using a GAS, which we can build offline and reuse since we have a static obstacle scene. We use the broad-phase collision detection results to collect potential colliding rays along transformed robot links. We observed it is better to transform link rays online rather than offline as it improves cache performance and thereby reduces global memory accesses. As with ObsToRobot, was also found 4 streams and 1024 poses per batch is the best configuration.
A-B3 CCD
Robot swept volumes change with each trajectory, preventing us from using IASs of pre-built GASs or performing fast dynamic updates of GASs. Instead, we are forced to build a new GAS for each batch of trajectories. We did try a custom curve segment OBB broad-phase to reduce the number of rays traced. However, this did not provide improvement, due to a combination of most of the space being filled with swept volumes and OBBs often being an inefficient bounding volume for swept sphere curves. Thus, the ray-tracing computation saved was outweighed by the extra overhead of running broad-phase collision detection. For this algorithm, we also landed on 4 streams but chose a smaller batch size of 128 trajectories per stream.
A-C cuRobo Benchmark Parameters
In our benchmark experiments, we made slight modifications to the cuRobo library and adjusted inputs to ensure a fair comparison. We adapted our cuRobo benchmark code from the benchmark code provided with the cuRobo library.
First, since our algorithms currently do not include self-collision detection, we disabled cuRobo’s self-collision detection. We also replaced cuRobo’s sphere representation of the Franka Panda robot with the same one used for our own algorithm. This sphere set has a few more spheres with less conservatism (shown in Fig. 2 and 6). For the CCD experiments, we set the cuRobo trajectory horizon to 8 and 16 to match our number of curve spline control points. We also set cuRobo’s sweep_step to 3, which gives it enough opportunity to capture collisions in dense obstacle environments.
We enabled CUDA graphs in all benchmarks to maximize cuRobo’s performance, resulting in a slight speedup and different trends across batch sizes. For comparison, the results with CUDA graphs disabled in cuRobo are shown in Fig. 7 and 8.
A-D CCD Strongly-Connected Directed Graph Construction
As explained in Sec. III-C, to avoid tracing two rays along each obstacle mesh edge in CCD, we build a strongly-connected directed graph on the each obstacle mesh’s edges and trace only along these directed graph edges, significantly reducing the number of rays required without sacrificing correctness. Due to limited space in the main paper, the explanation of our simple approach to construct this graph was quite brief, so we provide the pseudocode in Alg. 2.
Algorithm 2 Constructing a Strongly-Connected Directed Graph on Connected Triangle Mesh Edges
Input: Mesh triangles $T$, edge-to-triangle map
Output: Mesh edge directions
[…]
5:while $T_{\text{wavefront}}\neq\emptyset$ do
6:  $t\leftarrow$ Sample and remove triangle from $T_{\text{wavefront}}$
7:  $T_{\text{neighbors}}\leftarrow$ neighbors of $t$ in $T_{\text{processed}}$
8:  $n\leftarrow|T_{\text{neighbors}}|\;\;\blacktriangleright$ # of neighbor triangles (up to 3)
9:  if $n=0$ or $T_{\text{neighbors}}$ orientations are tied then
[…]
11:  else if all $T_{\text{neighbors}}$ orientations are the same then
12:   Set $t$ to opposite orientation $\quad\blacktriangleright n=1,2$, or 33
13:  else
[…]
15:  $T_{\text{processed}}\leftarrow T_{\text{processed}}\bigcup\{t\}$
16:  Move neighbors of $t$ in $T_{\text{unprocessed}}$ to $T_{\text{wavefront}}$
17:for edge in mesh do
18:  for triangle along edge do
19:   Add edge direction matching triangle orientation
20:return edge directions